OpenStudy (shamil98):

How to solve for x? 5 = log_3 (x^2 + 18)

OpenStudy (solomonzelman):

\[5 = \log_3 (x^2 + 18) \] let x^2 + 18=a so \[5 = log_3 a \] Does this help?

OpenStudy (shamil98):

Not really , my knowledge with logarithms is little... Dumb in this area, you could say.

OpenStudy (shamil98):

log_a x = N A^n = x so.. log_3 a = 5 3^5 = a ?

OpenStudy (solomonzelman):

\[\log_3(a) = 5\]\[Log_ab=c~~~~~~~->~~~~~~a^c=b~~~~~so,~~~~YES,~~you~~got~~it!\]

OpenStudy (shamil98):

Awesome! and the rest is algebra. 243 = x^2 + 18 x^2 - 225 (x+15)(x-15) x = 15, -15,

OpenStudy (solomonzelman):

243 = x^2 + 18 243 -18= x^2 and the answer is there CORRECT! 225=x^2 x=± 15

OpenStudy (shamil98):

thanks.

OpenStudy (solomonzelman):

Anytime!